OpenStudy (anonymous):

A certain first-order reaction has a rate constant of 2.80×10−2s−1 at 15∘C. What is the value of k at 62∘C if Ea = 112kJ/mol ?

OpenStudy (aaronq):

use the two-point arrhenius equation: \(ln\dfrac{k_1}{k_2}=\dfrac{E_a}{R}*(\dfrac{1}{T_2}-\dfrac{1}{T_1})\)

OpenStudy (anonymous):

I have tried that and it doesnt get the right answer :(

OpenStudy (aaronq):

you used Kelvin?

OpenStudy (anonymous):

yes and i changed kj to J

OpenStudy (aaronq):

let me see what i get

OpenStudy (aaronq):

are you getting 19.82?

OpenStudy (anonymous):

no -6.5

OpenStudy (anonymous):

do you have to do something with the k1 too?

OpenStudy (aaronq):

yeah, you have to fill all the parameters except what you're looking for

OpenStudy (anonymous):

cause if you do then i got .00004

OpenStudy (aaronq):

hm i got 19.82

OpenStudy (anonymous):

so 112000/8.314 * ((1/335)-(1/288)) = -6.56

OpenStudy (aaronq):

yeah, then \(ln(\dfrac{2.80×10^{−2}}{k_2})=-6.56\)

OpenStudy (anonymous):

oh i see you take the reciprocal ok i got your answer now

OpenStudy (aaronq):

sweet, that should be correct

OpenStudy (anonymous):

but the answer is supossed to be 2.3

OpenStudy (aaronq):

really? thats odd. is this a reliable source?

OpenStudy (anonymous):

mastering chemistry...online text software

OpenStudy (anonymous):

oh no we were right!! I just answered the wrong question lol. Thanks a lot it is right!!

OpenStudy (aaronq):

ohh ok, cool cool

OpenStudy (aaronq):

no prob